OpenStudy (iwanttogotostanford):

HELP ASAP:

Directrix (directrix):

First, solve x = 1/2 * t for t. t = ?

Directrix (directrix):

Post what you get, okay?

OpenStudy (iwanttogotostanford):

can you show me how

OpenStudy (518nad):

you might wanna loook how up algebra on youtube

OpenStudy (518nad):

t/2 = x solve for t

OpenStudy (518nad):

what happens if u multiply t/2 by 2

Directrix (directrix):

Multiply both sides by 2: x = 1/2 * t 2*x = 2 * 1/2 * t 2 * 1/2 = ?

Directrix (directrix):

I would like to help you with this but please help me help you by participating. Thanks.

OpenStudy (iwanttogotostanford):

is it A right?

OpenStudy (iwanttogotostanford):

@3mar i need math help plz

OpenStudy (3mar):

Well, I am here.

OpenStudy (iwanttogotostanford):

ok i posted the question please help ASAP

OpenStudy (3mar):

But are you going to waste my time waiting you?

OpenStudy (iwanttogotostanford):

no I'm here

OpenStudy (3mar):

Let's see :\

OpenStudy (3mar):

ques at the top?

OpenStudy (iwanttogotostanford):

yes

OpenStudy (3mar):

Ok

OpenStudy (iwanttogotostanford):

i just need the answer please no explain

OpenStudy (3mar):

Why?

OpenStudy (iwanttogotostanford):

because please just trust my process

OpenStudy (3mar):

Check you mean!

OpenStudy (iwanttogotostanford):

yes

OpenStudy (3mar):

Ok, as you did it, what is yours and I tell you correct or not!

OpenStudy (iwanttogotostanford):

ok i believe it is C or D am i right? please quick i have 10 more questions i really need help on

OpenStudy (3mar):

not correct sorry The answer is B

OpenStudy (3mar):

Need why?

OpenStudy (iwanttogotostanford):

ok thanks I need help please couple other questions can you help please?

OpenStudy (3mar):

Got that one?

OpenStudy (iwanttogotostanford):

yes i got that one thanks so much/

OpenStudy (iwanttogotostanford):

I believe this one is D

OpenStudy (3mar):

It is not useful to just give you the answer without you choose it and know why!

OpenStudy (iwanttogotostanford):

i know why please just help i know please just trust...

OpenStudy (3mar):

Can we start in new post, please?

OpenStudy (iwanttogotostanford):

ok